# Deformation of Fuchsian representations and proper affine actions

### v1

@@ Abstract.
The main goal of this article is to generalize Mess’ work and using results from Labourie–Wentworth and Potrie–Sambarino, to show that inside Hitchin representations, infinitesimal deformations of Fuchsian representations of a cocompact surface group do not act properly along the directions corresponding to $k$-differentials. In the process, we introduce affine versions of cross ratios and triple ratios. We introduce Margulis invariants and relate them with affine crossratios and infinitesimal Jordan projections. We obtain a general equivalent criterion for existence of proper affine actions in terms of the structure of the Margulis invariant spectra. Also, using a stability argument we show the existence of proper affine actions of non-abelian free groups whose linear part is a Hitchin representation.
@@ 1.1. Background and main results
Subsequently, Abels–Margulis–Soifer [AMS02] constructed free subgroups of $\mathsf{SO}_{2n,2n-1}\ltimes\mathbb{R}^{4n-1}$ which act properly on $\mathbb{R}^{4n-1}$. Moreover, they showed that there does not exist any free subgroup of $\mathsf{SO}_{2n+1,2n}\ltimes\mathbb{R}^{4n+1}$ which act properly on $\mathbb{R}^{4n+1}$. These results can be seen as generalizations of the first reformulation of Margulis’ work. Recently, generalizations of the second reformulation of Margulis’ work was obtained by Smilga [Smi16, Smi18, Smi22] and Danciger–Guéritaud–Kassel [DGK20]. When $\mathsf{G}$ is a noncompact semisimple real Lie group, Smilga [Smi16] constructed free subgroups of $\mathsf{G}\ltimes\mathfrak{g}$ which act properly on $\mathfrak{g}$. These results were further generalized by Smilga [Smi18, Smi22] to construct free subgroups of $\mathsf{G}\ltimes_{\mathtt{R}}\mathsf{V}$ which act properly on $\mathsf{V}$, under some conditions on the representation $\mathtt{R}$ via which the noncompact semsimple Lie group $\mathsf{G}$ act on $\mathsf{V}$. In particular the respresentation must admit zero as a weight. The linear parts of all these examples contain no parabolic elements, although Drumm [Dru93] showed that one could construct examples of free subgroups of $\mathsf{SO}_{2,1}\ltimes\mathbb{R}^{3}$ which act properly on $\mathbb{R}^{3}$ and whose linear part contain parabolic elements. All these constructions relied on a suitable generalization of the Margulis invariants. In general, roughly speaking the Margulis invariant of an element $(g,Y)\in\mathsf{G}\ltimes_{\mathtt{R}}\mathsf{V}$ is the standardized projection of $Y$ on the unit eigenspace of $g$ (see Definition 4.2.1 for more details). In this article, we introduce affine versions of cross ratios and relate them with Margulis invariants. Suppose $(g,Y)\in\mathsf{G}\ltimes_{\mathtt{R}}\mathsf{V}$ be such that its action on the space of affine parabolic subspaces (see Definitions 2.2.3 and 3.2.2) has an attracting fixed point $\mathsf{A}_{+}$, a repelling fixed point $\mathsf{A}_{-}$ and $\mathsf{A}_{\pm}$ are transverse to each other. Moreover, suppose $\mathsf{A}$ is another affine parabolic space which is transverse to both $A_{\pm}$ and $(g,Y)\mathsf{A}$ is transverse to all $\mathsf{A}$. We denote the affine cross ratio of the four mutually transverse affine parabolic subspaces by $\beta(\mathsf{A}_{+},\mathsf{A}_{-},(g,Y)\mathsf{A},\mathsf{A})$ (see Definition 4.1.4 for more details). We generalize results from Ghosh [Gho18] and prove that:
@@
In fact, we also introduce affine triple ratios and relate them with affine cross ratios (see Definition 4.1.6 and Proposition 4.1.7). Similar results relating Margulis invariants and cross ratios has also been independently obtained by Andrés Sambarino. Recently, there has been a lot of activity trying to characterize existence of proper affine actions in terms of the Margulis invariant spectra. These works show that examples whose linear part contain no parabolic elements can be characterized more easily than the examples constructed by Drumm. In particular, Goldman–Labourie–Margulis [GLM09] gave an equivalent criterion (see also Danciger–Guéritaud–Kassel [DGK16]), in terms of Margulis invariants, for existence of proper affine actions of free groups inside $\mathsf{SO}_{2,1}\ltimes\mathbb{R}^{3}$ whose linear part does not contain parabolic elements. Building upon this and the work of Goldman–Labourie [GL12], Ghosh [Gho17] related the equivalent criterion of proper affine actions with the notion of an Anosov representation. Anosov representations were introduced by Labourie [Lab06] to give a characterization of representations in the Hitchin components [Hit92]. The most important property of Anosov representations is that they are stable under small deformations. Let $\Gamma$ be a word hyperbolic group, $\tau:\Gamma\to\mathsf{SL}_{2}(\mathbb{R})$ be any injective homomorphism and $\iota:\mathsf{SL}_{2}(\mathbb{R})\to\mathsf{G}$ be the irreducible representation. Then any representation of the form $\iota\circ\tau$ is called a Fuchsian representation and the connected components of the representation variety which contain Fuchsian representations are called Hitchin components. We denote these components by $\mathsf{Hit}(\Gamma,\mathsf{G})$. Any representation contained in the Hitchin components is called a Hitchin representation. Subsequently, Ghosh–Trieb [GT17] generalized the result of Goldman–Labourie–Margulis and gave an equivalent criterion, in terms of Margulis invariants and Anosov representations, for existence of proper affine actions of word hyperbolic groups inside $\mathsf{SO}_{2n,2n-1}\ltimes\mathbb{R}^{4n-1}$ whose linear part does not contain parabolic elements. We extend these results in this paper by replacing $\mathsf{SO}_{2n,2n-1}\ltimes\mathbb{R}^{4n-1}$ with $\mathsf{G}\ltimes_{\mathtt{R}}\mathsf{V}$, where $\mathsf{G}$ is a real split semisimple algebraic Lie group with trivial center and $\mathtt{R}:\mathsf{G}\to\mathsf{SL}(\mathsf{V})$ is a faithful irreducible algebraic non-swinging representation admitting zero as a weight (see Remark 2.1.3).
@@
We note that Kassel–Smilga [KSon] have a similar theorem in a related, but slightly different, setting. Moreover, when $\mathsf{V}=\mathfrak{g}$ and $\mathtt{R}$ is the adjoint representation we generalize results of Goldman–Margulis [GM00] and show in Proposition 4.2.4 that the Margulis invariants are infinitesimal Jordan projections. Similar results relating Margulis invariants and Jordan projections has also been independently obtained by Andrés Sambarino and Kassel–Smilga [KSon]. Furthermore, we use the stability of Anosov representations under small deformations to prove the following existence result (similar result but in a different setting was previously obtained by Ghosh [Gho18]):
@@
On contrary, Mess [Mes07] showed that the fundamental group of a closed compact surface without boundary admits no proper affine action on $\mathbb{R}^{3}$. Later, Goldman–Margulis [GM00] gave an alternate proof of Mess’ result. This result can be reformulated in the following two ways:
@@
    Suppose $\Gamma$ is the fundamental group of a closed compact surface without boundary and $(\rho,u):\Gamma\to\mathsf{SO}_{2,1}\ltimes\mathbb{R}^{3}$ be an injective homomorphism such that $\rho$ is Fuchsian. Then $(\rho,u)(\Gamma)$ does not act properly on $\mathbb{R}^{3}$.
@@
    Suppose $\Gamma$ is the fundamental group of a closed compact surface without boundary and $(\rho,u):\Gamma\to\mathsf{SL}_{2}(\mathbb{R})\ltimes\mathfrak{sl}_{2}(\mathbb{R})$ be an injective homomorphism. Then $(\rho,u)(\Gamma)$ does not act properly on $\mathfrak{sl}_{2}(\mathbb{R})$.
Subsequently, Labourie [Lab01] showed that if $\Gamma$ is the fundamental group of a closed compact surface without boundary and $(\rho,u):\Gamma\to\mathsf{SO}_{n+1,n}\ltimes\mathbb{R}^{2n+1}$ be such that $\rho$ is Fuchsian, then $(\rho,u)(\Gamma)$ does not act properly on $\mathbb{R}^{2n+1}$. Recently, Danciger–Zhang [DZ19] and Labourie [Lab22] generalized Labourie’s result and showed that if $\Gamma$ is the fundamental group of a closed compact surface without boundary and $(\rho,u):\Gamma\to\mathsf{SO}_{n+1,n}\ltimes\mathbb{R}^{2n+1}$ be such that $\rho$ is Hitchin, then $(\rho,u)(\Gamma)$ does not act properly on $\mathbb{R}^{2n+1}$. Both these results generalize the first reformulation of Mess’ result. In this article, we generalize the second reformulation of Mess’ result and show the following:
@@
Let $\Gamma$ be a cocompact surface group and $(\rho,u):\Gamma\to\mathsf{SL}_{n}(\mathbb{R})\ltimes\mathfrak{sl}_{n}(\mathbb{R})$ is an injective homomorphism such that $\rho$ is a Fuchsian representation. Also, suppose $[u]$ in $\mathsf{T}_{[\rho]}\mathsf{Hit}(\Gamma,\mathsf{SL}_{n}(\mathbb{R}))$ correspond to a $k$-differential. Then $(\rho,u)(\Gamma)$ does not act properly on $\mathfrak{sl}_{n}(\mathbb{R})$.
As a main ingredient in proving the above result we use results of Labourie–Wentworth [LW18] and Potrie–Sambarino [PS17]. More generally, we are interested in knowing the answer to the following question:
@@
The naive guess would be that $(\rho,u)(\Gamma)$ does not act properly on $\mathfrak{sl}_{n}(\mathbb{R})$ for all $[u]$ in $\mathsf{T}_{[\rho]}\mathsf{Hit}(\Gamma,\mathsf{SL}_{n}(\mathbb{R}))$ but we are unable to prove this result. An immediate reason is due to the existing restrictions in the statement of Corollary 1.4 of Potrie–Sambarino [PS17] but a more major reason is the fact that the Margulis invariant spectra is not in general contained in a one dimensional subspace.
@@ 1.2. Plan of the paper
The paper consists of four main sections. In the first section we introduce preliminary notions which are needed later to state and prove our results. The first section is divided into two subsections. In subsection 2.1 we recall some known results about representations of Lie groups. In subsection 2.2 we define Anosov representations in general and state some of its properties relevant to our article. In the second section we study affine deformations of Anosov representations. This section is again divided into two parts. In subsection 3.1 we introduce neutral sections, a concept first introduced by Goldman–Labourie–Margulis [GLM09] to set the ground of defining affine equivalent of an Anosov representation. And, in subsection 3.2 we define appropriate affine versions of Anosov representations. In the third section we introduce certain affine invariants which play a central role in our article. The third section is divided into three subsections. In subsection 4.1 we introduce affine cross ratios and affine tripe ratios and show their inter-relationship. In subsection 4.2 we introduce Margulis invariants and relate them with affine cross ratios and infinitesimal Jordan projections. In subsection 4.3 we prove a limit formula and obtain convexity results about the Margulis invariant spectra and the Jordan-Margulis invariant spectra. In the fourth section we prove results related to proper affine actions. This section is divided into three parts. In subsection 5.1 we prove one of our main results and give a general criterion about proper affine actions and derive some other useful criteria about proper affine actions from it. In subsection 5.2 we use the stability of Anosov representations and show existence of proper affine actions of free surface groups with Hitchin linear part. Finally, in subsection 5.3 we prove the most important result of this article. We use results from [LW18] and [PS17] to obtain non-existence results of proper affine actions of cocompact surface groups with Fuchsian linear part.
@@ 1.3. Acknowledgements
I would like to thank François Labourie for his insight, encouragement and in particular for funding me to visit him at Nice. I would also like to thank him for his suggestions to improve the exposition of this article and C. S. Rajan for many fruitful discussions. Moreover, I would like to heartily thank all those people who were beside me during the last couple of years. In particular, my parents Swapan Ghosh, Purnima Ghosh, my brother Gourab Ghosh, my well wishers Sandeepan Parekh, Nishant Chandgotia, Arghya Mondal, Neeraja Sahasrabudhe, Pallavi Panda, Indira Chatterji, François Labourie, Mahan Mj, Indranil Biswas, Jean-Marc Schlenker, Anna Wienhard, Cagri Sert and Rajendra Bhatia. Without their timely support this article could never have been possible. Lastly, we acknowledge that Andrés Sambarino, Fanny Kassel and Ilia Smilga have independently derived results similar to some of the results obtained in this article.

### v3

@@ Abstract.
The main goal of this article is to generalize Mess’ work and using results from Labourie–Wentworth, Potrie–Sambarino and Smilga, to show that inside Hitchin representations, infinitesimal deformations of Fuchsian representations of a cocompact surface group do not act properly along the directions corresponding to the sum of a mixed odd differential and a $2m$-differential for any $1\leq m\leq\lfloor\frac{n}{2}\rfloor$.
In the process, we introduce affine versions of cross ratios and triple ratios. We introduce Margulis invariants and relate them with affine crossratios and infinitesimal Jordan projections. We obtain a general equivalent criterion for existence of proper affine actions in terms of the structure of the Margulis invariant spectra. Also, using a stability argument we show the existence of proper affine actions of non-abelian free groups whose linear part is a Hitchin representation.
@@ 1.1. Background and main results
Subsequently, Abels–Margulis–Soifer [AMS02] constructed free subgroups of $\mathsf{SO}_{2n,2n-1}\ltimes\mathbb{R}^{4n-1}$ which act properly on $\mathbb{R}^{4n-1}$. Moreover, they showed that there does not exist any free subgroup of $\mathsf{SO}_{2n+1,2n}\ltimes\mathbb{R}^{4n+1}$ which act properly on $\mathbb{R}^{4n+1}$. These results can be seen as generalizations of the first reformulation of Margulis’ work. Recently, generalizations of the second reformulation of Margulis’ work was obtained by Smilga [Smi16, Smi18, Smi22] and Danciger–Guéritaud–Kassel [DGK20]. When $\mathsf{G}$ is a noncompact semisimple real Lie group, Smilga [Smi16] constructed free subgroups of $\mathsf{G}\ltimes\mathfrak{g}$ which act properly on $\mathfrak{g}$. These results were further generalized by Smilga [Smi18, Smi22] to construct free subgroups of $\mathsf{G}\ltimes_{\mathtt{R}}\mathsf{V}$ which act properly on $\mathsf{V}$, under some conditions on the representation $\mathtt{R}$ via which the noncompact semsimple Lie group $\mathsf{G}$ acts on $\mathsf{V}$. In particular the respresentation must admit zero as a weight. The linear parts of all these examples contain no parabolic elements, although Drumm [Dru93] showed that one could construct examples of free subgroups of $\mathsf{SO}_{2,1}\ltimes\mathbb{R}^{3}$ which act properly on $\mathbb{R}^{3}$ and whose linear part contains parabolic elements. All these constructions relied on a suitable generalization of the Margulis invariants. In general, roughly speaking the Margulis invariant of an element $(g,Y)\in\mathsf{G}\ltimes_{\mathtt{R}}\mathsf{V}$ is the standardized projection of $Y$ on the unit eigenspace of $g$ (see Definition 4.2.1 for more details). In this article, we introduce affine versions of cross ratios and relate them with Margulis invariants. Suppose $(g,Y)\in\mathsf{G}\ltimes_{\mathtt{R}}\mathsf{V}$ be such that its action on the space of affine parabolic subspaces (see Definitions 2.2.3 and 3.2.2) has an attracting fixed point $\mathsf{A}_{+}$, a repelling fixed point $\mathsf{A}_{-}$ and $\mathsf{A}_{\pm}$ are transverse to each other. Moreover, suppose $\mathsf{A}$ is another affine parabolic space which is transverse to both $A_{\pm}$ and $(g,Y)\mathsf{A}$ is transverse to $\mathsf{A}$. We denote the affine cross ratio of the four mutually transverse affine parabolic subspaces by $\beta(\mathsf{A}_{+},\mathsf{A}_{-},(g,Y)\mathsf{A},\mathsf{A})$ (see Definition 4.1.4 for more details). We generalize results from Ghosh [Gho18] and prove that:
@@
In fact, we also introduce affine triple ratios and relate them with affine cross ratios (see Definition 4.1.6 and Proposition 4.1.7). Similar results relating Margulis invariants and cross ratios has also been independently obtained by Andrés Sambarino [Sam24].
Recently, there has been a lot of activity trying to characterize existence of proper affine actions in terms of the Margulis invariant spectra. These works show that examples whose linear part contain no parabolic elements can be characterized more easily than the examples constructed by Drumm. In particular, Goldman–Labourie–Margulis [GLM09] gave an equivalent criterion (see also Danciger–Guéritaud–Kassel [DGK16]), in terms of Margulis invariants, for existence of proper affine actions of free groups inside $\mathsf{SO}_{2,1}\ltimes\mathbb{R}^{3}$ whose linear part does not contain parabolic elements. Building upon this and the work of Goldman–Labourie [GL12], Ghosh [Gho17] related the equivalent criterion of proper affine actions with the notion of an Anosov representation. Anosov representations were introduced by Labourie [Lab06] to give a characterization of representations in the Hitchin components [Hit92]. A fundamental property of Anosov representations is that they are stable under small deformations. Let $\Gamma$ be a word hyperbolic group, $\tau:\Gamma\to\mathsf{SL}_{2}(\mathbb{R})$ be any injective homomorphism and $\iota:\mathsf{SL}_{2}(\mathbb{R})\to\mathsf{G}$ be the irreducible representation. Then any representation of the form $\iota\circ\tau$ is called a Fuchsian representation and the connected components of the representation variety which contain Fuchsian representations are called Hitchin components. We denote these components by $\mathsf{Hit}(\Gamma,\mathsf{G})$. Any representation contained in the Hitchin components is called a Hitchin representation. Subsequently, Ghosh–Treib [GT17] generalized the result of Goldman–Labourie–Margulis and gave an equivalent criterion, in terms of Margulis invariants and Anosov representations, for existence of proper affine actions of word hyperbolic groups inside $\mathsf{SO}_{2n,2n-1}\ltimes\mathbb{R}^{4n-1}$ whose linear part does not contain parabolic elements. We extend these results in this paper by replacing $\mathsf{SO}_{2n,2n-1}\ltimes\mathbb{R}^{4n-1}$ with $\mathsf{G}\ltimes_{\mathtt{R}}\mathsf{V}$, where $\mathsf{G}$ is a real split semisimple algebraic Lie group with trivial center and $\mathtt{R}:\mathsf{G}\to\mathsf{SL}(\mathsf{V})$ is a faithful irreducible algebraic non-swinging representation admitting zero as a weight (see Remark 2.1.3).
@@
We note that Kassel–Smilga [KSon] have a similar theorem in a related, but slightly different, setting.
Moreover, when $\mathsf{V}=\mathfrak{g}$ and $\mathtt{R}$ is the adjoint representation we generalize results of Goldman–Margulis [GM00] and show in Proposition 4.2.4 that the Margulis invariants are infinitesimal Jordan projections. Similar results relating Margulis invariants and Jordan projections has also been independently obtained by Andrés Sambarino [Sam24] and Kassel–Smilga [KSon]. Furthermore, we use the stability of Anosov representations under small deformations to prove the following existence result (similar result but in a different setting was previously obtained by Ghosh [Gho18]):
@@
On contrary, Mess [Mes07] showed that the fundamental group of a compact orientable surface without boundary and of genus atleast two admits no proper affine action on $\mathbb{R}^{3}$. Later, Goldman–Margulis [GM00] gave an alternate proof of Mess’ result. This result can be reformulated in the following two ways:
@@
    Suppose $\Gamma$ is the fundamental group of a compact orientable surface without boundary and of genus atleast two and $(\rho,u):\Gamma\to\mathsf{SO}_{2,1}\ltimes\mathbb{R}^{3}$ is an injective homomorphism such that $\rho$ is Fuchsian. Then $(\rho,u)(\Gamma)$ does not act properly on $\mathbb{R}^{3}$.
@@
    Suppose $\Gamma$ is the fundamental group of a compact orientable surface without boundary and of genus atleast two and $(\rho,u):\Gamma\to\mathsf{SL}_{2}(\mathbb{R})\ltimes\mathfrak{sl}_{2}(\mathbb{R})$ be an injective homomorphism. Then $(\rho,u)(\Gamma)$ does not act properly on $\mathfrak{sl}_{2}(\mathbb{R})$.
Subsequently, Labourie [Lab01] showed that if $\Gamma$ is the fundamental group of a compact orientable surface without boundary and of genus atleast two and $(\rho,u):\Gamma\to\mathsf{SO}_{n+1,n}\ltimes\mathbb{R}^{2n+1}$ is such that $\rho$ is Fuchsian, then $(\rho,u)(\Gamma)$ does not act properly on $\mathbb{R}^{2n+1}$. Recently, Danciger–Zhang [DZ19] and Labourie [Lab22] generalized Labourie’s result and showed that if $\Gamma$ is the fundamental group of a compact orientable surface without boundary and of genus atleast two and $(\rho,u):\Gamma\to\mathsf{SO}_{n+1,n}\ltimes\mathbb{R}^{2n+1}$ is such that $\rho$ is Hitchin, then $(\rho,u)(\Gamma)$ does not act properly on $\mathbb{R}^{2n+1}$. Both these results generalize the first reformulation of Mess’ result. In this article, we generalize the second reformulation of Mess’ result and show the following:
@@
Let $\Gamma$ be a cocompact surface group and $(\rho,u):\Gamma\to\mathsf{SL}_{n}(\mathbb{R})\ltimes\mathfrak{sl}_{n}(\mathbb{R})$ is an injective homomorphism such that $\rho$ is a Fuchsian representation. Also, suppose $[u]$ in $\mathsf{T}_{[\rho]}\mathsf{Hit}(\Gamma,\mathsf{SL}_{n}(\mathbb{R}))$ correspond to the sum of a mixed odd differential and a $2m$-differential for any $1\leq m\leq\lfloor\frac{n}{2}\rfloor$. Then $(\rho,u)(\Gamma)$ does not act properly on $\mathfrak{sl}_{n}(\mathbb{R})$.
As a main ingredient in proving the above result we use results of Smilga [Smi22], Labourie–Wentworth [LW18] and Potrie–Sambarino [PS17]. More generally, we are interested in knowing the answer to the following question:
@@
The naive guess would be that $(\rho,u)(\Gamma)$ does not act properly on $\mathfrak{sl}_{n}(\mathbb{R})$ for all $[u]$ in $\mathsf{T}_{[\rho]}\mathsf{Hit}(\Gamma,\mathsf{SL}_{n}(\mathbb{R}))$ but we are unable to prove this result. An immediate reason is due to the existing restrictions in the statement of Corollary 1.4 of Potrie–Sambarino [PS17] but a more major reason is the fact that no relation is known to exist (at least by us) between Margulis invariant of a single element along different $k$-differentials.
@@ 1.2. Plan of the paper
The paper consists of four main sections.
In the first section we introduce preliminary notions which are needed later to state and prove our results. The first section is divided into two subsections. In subsection 2.1 we recall some known results about representations of Lie groups. In subsection 2.2 we define Anosov representations in general and state some of their properties relevant to our article.
In the second section we study affine deformations of Anosov representations. This section is again divided into two parts. In subsection 3.1 we introduce neutral sections, a concept first introduced by Goldman–Labourie–Margulis [GLM09] to set the ground of defining affine equivalent of an Anosov representation. And, in subsection 3.2 we define appropriate affine versions of Anosov representations.
In the third section we introduce certain affine invariants which play a central role in our article. The third section is divided into three subsections. In subsection 4.1 we introduce affine cross ratios and affine tripe ratios and show their inter-relationship. In subsection 4.2 we introduce Margulis invariants and relate them with affine cross ratios and infinitesimal Jordan projections. In subsection 4.3 we prove a limit formula and obtain convexity results about the Margulis invariant spectra and the Jordan-Margulis invariant spectra.
In the fourth section we prove results related to proper affine actions. This section is divided into three parts. In subsection 5.1 we prove one of our main results and give a general criterion about proper affine actions and derive some other useful criteria about proper affine actions from it. In subsection 5.2 we use the stability of Anosov representations and show existence of proper affine actions of free surface groups with Hitchin linear part. Finally, in subsection 5.3 we prove the most important result of this article. We use results from [LW18], [PS17] and Smilga [Smi22] to obtain non-existence results of proper affine actions of cocompact surface groups with Fuchsian linear part.
@@ 1.3. Acknowledgements
I would like to thank François Labourie for his insight, encouragement and in particular for funding me to visit him at Nice. I would also like to thank him for his suggestions to improve the exposition of this article and Pritam Ghosh and C. S. Rajan for many fruitful discussions. Moreover, I would like to heartily thank all those people who were beside me during the last couple of years. In particular, my parents Swapan Ghosh, Purnima Ghosh, my brother Gourab Ghosh, my well wishers Sandeepan Parekh, Nishant Chandgotia, Arghya Mondal, Neeraja Sahasrabudhe, Pallavi Panda, Indira Chatterji, François Labourie, Mahan Mj, Indranil Biswas, Jean-Marc Schlenker, Anna Wienhard, Cagri Sert and Rajendra Bhatia. Without their timely support this article could never have been possible.
Lastly, we acknowledge that Andrés Sambarino, Fanny Kassel and Ilia Smilga have independently derived results similar to some of the results obtained in this article.